# DeepScholar-Bench: A Live Benchmark and Automated Evaluation for Generative Research Synthesis

### v1

@@ 3.2. Retrieval Quality
While traditional information retrieval (IR) evaluations typically leverage gold labels for document relevance scores and a controlled corpus (Thakur et al., 2021; noa, [n.d.]l; Nanni et al., 2017), the research synthesis task we study in this work differs substantially. An expert-written report section and its sources provide one reasonable retrieved set, but there may be many possible alternative sets that are likewise high-quality. Moreover, live web search is a key component of research synthesis tasks and obtaining gold relevance labels over this corpus is prohibitively expensive. To address these challenges, our evaluation measures three components of the retrieved set: the relevance rate, reference coverage of key sources, and document importance.
Relevance Rate. We asses the relevance of each retrieved document, following the Cranfield model (Voorhees, 2009), which is standard in IR evaluations and considers relevance of individual documents given a query, independent of other documents. Due to the significant cost of obtaining human-annotated relevance judgments, recent works (Upadhyay et al., 2024b; Faggioli et al., 2023; Rahmani et al., 2024b; Thomas et al., 2024; Asai et al., 2024) study leveraging an LLM-as-a-judge for relevance judgment task, demonstrating their effectiveness on traditional IR datasets. Building on these works, we adopt an LLM-based approach for assigning graded relevance scores from 0 to 2 to each generated research report using the prompt in Box 8.4.3 in Appendix. For each retrieved set, , for a given query, we compute the average document relevance over the set, following the below formula:
where is the graded relevance score of source . We validate the agreement between LLMs and human annotators for this task in our experiments in Section 5.
@@
Reference Coverage. We introduce a metric to measure the reference coverage of the retrieved set for each report. A key challenge in measuring this value is in defining a set of important references for each generated report, that a good research report should cite. To build this set, we take all references from the high-quality, human-written exemplar and label each as either ”important” or ”not-important”, considering a ”not-important” reference as one that could be omitted or substituted by a different reference. We find that a LLM-based judge is effective in assigning these labels. For a given report, we then compute its reference coverage by taking the ratio of the number of important references in the system-generated report to the number of important references in the human-written exemplar, following the below formula, where is the set of ”important” references from the human-written exemplar:
@@
Document Importance. While the above relevance and coverage metrics assess topical matches between the retrieved set and the user query, an ideal research synthesis system must also retrieve notable and important sources. Exemplar human-written reports typically contain ample references of primary-sources and highly-cited academic publications. While the ideal notion of document importance depends on the particular task and user, in our task, we adopt the following metric by considering the number of citations that each reference retrieved by the system has. We consider the median number of citations per reference over the set of all retrieved sources, , provided by a given baseline. We compute document importance as the ratio of this value for the given baseline compared to the median citations per reference over the set of sources, provided by the human-written exemplars, following the formula below:
where is the number of citations for source . We set an upper-bound of one, although in practice, we find this ratio to remains far below one for all measured generative research synthesis systems.
3.3. Verifiability
@@
To evaluate the verifiability of the generated report given the retrieved set, we calculate the citation precision and claim coverage based on the definitions provided by prior work (Gao et al., 2023; Worledge et al., 2024; Liu et al., 2023).
@@
Citation Precision. Specifically, we measure sentence-level precision, where a citation is considered precise if the referenced source supports at least one claim made in the accompanying sentence. For a full report, citation precision is computed by averaging the precision of each citation in the report.
@@
Claim Coverage. Claim coverage assigns a sentence-level score, assigning a value of one if the cited sources accompanying the sentence supports all claims made in the sentence. We make two adaptations to the original definition posed in prior work (Gao et al., 2023; Worledge et al., 2024; Liu et al., 2023) for our long-form synthesis task. First, we relax the original claim coverage definition to consider a sliding window of sentences with supporting references, assigning a coverage value of 1 to each sentence that is either fully supported by the sources cited within the sentence or any cited source that is in a window of preceding or following sentences. Additionally, since our task query provides context describing a paper, we consider this context as an implicitly cited reference for each sentence and compute claim coverage for each sentence by considering the explicitly cited sources, and the context provided by the query. We compute the citation coverage for the full report by averaging the value computed for each sentence. Our model-based evaluation uses an LLM judge to assess each entailment relation, following prior work (Gao et al., 2023; Liu et al., 2023). We provide further details and the prompt in the Appendix Section 8.4 and Box 8.4.3 respectively.
4. DeepScholar-base
@@
We introduce DeepScholar-base, an open-source reference pipeline designed to perform generative research synthesis. Figure 3 provides an overview of our method. Given a user’s query, DeepScholar-base iteratively generates web-search queries, summarizes the search results in each round before generating a new query. The system then post-processes the search results leveraging a series of semantic operators (Patel et al., 2025), which we implement efficiently using the LOTUS API (noa, 2025b). This includes a semantic filtering step, which leverages an LLM to filter out irrelevant source documents, followed by a semantic top-k which performs an LLM-based ranking over the documents based on their relevance to the user query. Finally, we perform a semantic aggregation over the final source documents to generate the final report. We provide further details of each step of our reference pipeline in Appendix Section 8.3.
5. Experimental Results
@@
In this section, we evaluate recent state-of-the-art generative research systems as well as DeepScholar-base on DeepScholar-bench. Overall, we find the following:
@@
Existing baselines for generative research synthesis, including strong open-source LLM systems, search AI’s, and commercial systems, demonstrate significant room for improvement across all three key dimensions: knowledge synthesis, retrieval quality and verifiability, with no method exceeding a score of across all metrics.
@@
DeepScholar-base provides a strong baseline for generative research synthesis, consistently improving upon the performance of prior open-source systems and search AI’s, as well as achieving competitive performance and up to higher verifiability compared to OpenAI’s DeepResearch.
Our automated evaluation approach is effective, demonstrating high agreement with over 200 human annotations.

### v2

@@ 3.2 Retrieval Quality
A key component of generative research synthesis is retrieval over the live web, which differs substantially from traditional information retrieval evaluations [47, 32]. This setting lacks a closed corpus with gold labels — expert-written exemplars provide one reasonable reference set, but there may be many possible alternative sets that are likewise high-quality. To address these challenges, we evaluate three metrics of each generated report’s retrieved reference set: the relevance rate, reference coverage of key sources, and document importance.
Relevance Rate. We asses the relevance of each retrieved document, following the Cranfield model [52], which is standard in IR evaluations and considers relevance of individual documents given a query, independent of other documents. We use an LLM-as-a-judge approach to assign graded relevance scores to each retrieved source, following recent works [51, 9, 42, 48, 6]. Specifically, the LLM-judge assigns a relevance score, , from for each source, , in the retrieved set , and we compute the relevance rate of as:
@@
Reference Coverage. We introduce a metric to measure the reference coverage of each report’s retrieved set. A key challenge in measuring this value is in defining a core set of "important" sources that a good report should reference. To build this set, we take all references from the high-quality, human-written exemplar and label each as either "important" or "not-important", considering a "not-important" reference as one that could be omitted or substituted by a different source. For each generated report, we then report its reference coverage by dividing the number of retrieved important references by total number of important references. We follow the below formula, where is the set of "important" references from the human-written exemplar:
@@
Document Importance. While the above retrieval metrics assess topical relevance and coverage of key references, an ideal research synthesis system must also retrieve many notable and important sources. Exemplar human-written reports typically contain ample references of primary-sources and highly-cited academic publications. We compute the document importance of a retrieved set by considering the number of citations of each of its sources. Specifically, we consider the median number of citations over sources in , compared to this the median number of citations over sources in the reference set, , from the human-written exemplar, and set an upper-bound of 1, given by:
where is the number of citations for source .
3.3 Verifiability
@@
To evaluate the verifiability generated reports using the citation precision and claim coverage with LLM-based entailment evaluations, following prior work [11, 26].
@@
Citation Precision. We measure sentence-level precision, where a citation is considered precise if the referenced source supports at least one claim made in the accompanying sentence. For a full report, we compute citation precision by averaging the precision of each citation.
@@
Claim Coverage. Claim coverage of a report is computed by assigning a sentence-level score—of one if the sentence’s cited sources support all claims made in the sentence—and averaging all sentence-level scores in the report. We make two adaptations to the original definition of prior work [11, 57, 26] to tailor our metric to our long-form synthesis task. First, we relax the original claim coverage definition to consider a sliding window of sentences with supporting references Specifically, we assign a sentence-level coverage score of 1 if the sentence is fully supported by the sources cited either within the sentence or in a window of preceding or following sentences. Additionally, since our task query provides context describing a paper, we consider this context as an implicitly cited reference for each sentence.
4 DeepScholar-ref
@@
We introduce DeepScholar-ref, an open-source reference pipeline designed for generative research synthesis. As Figure 2 shows, DeepScholar-ref takes a user’s query and iteratively generates web-search queries, summarizing search results in each round before generating new queries. The system then post-processes the search results leveraging a series of semantic operators [37] implemented on the LOTUS API [28]. This includes a semantic filtering step, which leverages an LLM to filter out irrelevant source documents, followed by a semantic top-k which performs an LLM-based ranking over the documents based on their relevance to the user query. Finally, we perform a semantic aggregation over the final source documents to generate the final report. We provide further details of each step of our reference pipeline in Appendix Section A.6.
5 Experimental Results
@@
In this section, we evaluate recent state-of-the-art generative research systems as well as DeepScholar-ref on DeepScholar-bench. Overall, we find the following:
@@
Existing baselines for generative research synthesis, including strong open-source LLM systems, search agents, and commercial systems, demonstrate significant room for improvement across all three key dimensions: knowledge synthesis, retrieval quality and verifiability. Specifically, no system surpasses a geometric mean of across all metrics.
@@
DeepScholar-ref provides a strong baseline, consistently improving upon the performance of prior open-source systems and search agents, as well as achieving competitive performance and up to higher verifiability compared to OpenAI’s DeepResearch.
Experimental Setup. We benchmark open-source research systems, including DeepResearcher [65], STORM [43] and OpenScholar [6], search agents, with Llama-4-Scout-17B-16E-Instruct [30], GPT-4.1-2025-04-14 [34], o3-2025-04-16 [34], Claude-opus-4-20250514 [3], and Gemini-2.5-pro [14] models, OpenAI’s o3-deep-research [34], and DeepScholar-ref. For each benchmarked method, we control the retrieval corpus by allowing each system to access the Web only through the ArXiv API [5]. We additionally avoid possible information leakage during search by filtering out any search results that were published after the query paper’s publication date. We provide further details of our setup in Appendix A.2.
5.1 Main Results
Table 2 provides of each method’s performance on DeepScholar-June-2025. For each baseline, we report metrics averaged over all queries, and the geometric mean of all metrics. We also provide additional results in Appendix A.3—including metadata statistics of the generated reports (Table 6), statistics related to our evaluation metrics (Table 7), and manual validation of our LLM-based metrics (Table 10)—and example reports in Appendix A.7. We discuss key findings in detail below.
| Knowledge Synthesis | Retrieval Quality | Verifiability | Geo. Mean | |||||
| Org. | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov () | ||
| Human-written Exemplars | ||||||||
| Human-written Exemplars | .500 | 1.000 | .585 | 1.000 | 1.000 | .90011footnotemark: 1 | .85011footnotemark: 1 | .78211footnotemark: 1 |
| Open Source Research Systems | ||||||||
| DeepResearcher (Llama-4) | .206 | .230 | .385 | .047 | .008 | .312 | .396 | .137 |
| STORM (Llama-4) | .119 | .183 | .218 | .003 | .006 | .238 | .586 | .073 |
| OpenScholar (Llama-4) | .309 | .278 | .017 | .008 | .013 | .010 | .138 | .042 |
| Search Agents | ||||||||
| Search Agent (Llama-4) | .151 | .193 | .445 | .060 | .009 | .316 | .368 | .135 |
| Search Agent (GPT-4.1) | .556 | .265 | .490 | .050 | .009 | .498 | .470 | .186 |
| Search Agent (o3) | .849 | .348 | .610 | .165 | .026 | .425 | .495 | .287 |
| Search Agent (Claude) | .698 | .307 | .583 | .131 | .008 | .701 | .760 | .256 |
| Search Agent (Gemini) | .706 | .277 | .583 | .061 | .010 | .415 | .398 | .196 |
| Commercial Systems | ||||||||
| OpenAI DeepResearch | .857 | .392∗ | .629 | .187∗ | .124∗ | .399 | .138 | .309∗ |
| DeepScholar Reference Pipeline | ||||||||
| DeepScholar-ref (Llama-4) | .206 | .241 | .436 | .103 | .008 | .674 | .851 | .195 |
| DeepScholar-ref (GPT-4.1) | .809 | .348 | .590 | .166 | .008 | .788 | .899 | .285 |
| DeepScholar-ref (GPT-4.1, o3) | .857 | .384 | .645 | .167 | .007 | .824 | .760 | .285 |
| DeepScholar-ref (GPT-4.1, Claude) | .698 | .307 | .610 | .152 | .009 | .944∗ | .895 | .286 |
| DeepScholar-ref (GPT-4.1, Gemini) | .770 | .331 | .590 | .181 | .006 | .904 | .937∗ | .282 |
1
The automated verifiability metrics in our evaluation under-estimate the actual verifiability of human writing, thus, we provide an estimate using manual validation over a small sample, and we disclude them from the geometric mean for the human-written exemplars. This is because Citation Precision and Claim Coverage require us to assess entailment relations between claims and cited reference. For each LLM-based system, we are able to track the precise snippet and context from cited sources, which are directly fed as context to the LLM. On the other hand, for the human-written exemplars, we lack gold labels pointing to the precise snippet of text that each reference refers to. Our measurements for the human-written exemplars instead rely on the title and abstract of each cited source as a proxy.
5.1.1 Generative Research Synthesis Systems Demonstrate Large Room for Improvement.
From Table 2, we see that no system surpasses a geometric mean of .31 across all metrics, with OpenAI DeepResearch obtaining the highest geometric mean. Moreover, on several key metrics, including Nugget Coverage, Reference Coverage and Document Importance, each baseline’s performance remains below . This reflects the inherent difficulty of the generative research task provided by DeepScholar-bench, particularly the need for systems to navigate the live web, reasoning about coverage and importance of documents, before surfacing key information in long-form report.
We now analyze each evaluated dimension, comparing performance of the open-source research systems, search agents and commercial systems to the human-written exemplars. On Knowledge Synthesis, we see that OpenAI DeepResearch offers the best performance compared to all other prior methods on both Organization, with a score of , and Nugget Coverage, with a score of . OpenAI DeepResearch, as well as the o3, Claude and Gemini search agents achieve relatively high Organization scores compared to human-written exemplars. However, on Nugget Coverage all prior methods scores below . This demonstrates that while existing systems, especially those using state-of-the-art models, can generate well-organized and coherent summaries, they still struggle to surface key facts to answer the research query, a crucial capability for research synthesis tasks.
Turning our attention to the Retrieval Quality performance of prior methods, we once again find significant room for improvement. Once again, OpenAI DeepResearch offers the strongest performance among the other benchmarked prior methods on Relevance Rate, Reference Coverage and Document Importance, but still far from saturates performance. While it’s Relevance Rate shows strong performance, exceeding that of the human exemplars with a score of , it’s Reference Coverage and Document Importance scores remain exceedingly low: and respectively. This demonstrates that while state-of-the-art generative research synthesis systems excel at retrieving relevant sources, they still struggle to find comprehensive sets of notable sources, falling short compared to the ability of human experts.
Lastly, we analyze the Verifiability performance of prior methods. We see that OpenAI DeepResearch is outperformed on both Citation Precision and Claim Coverage by the search agents with GPT4.1, o3, Claude and Gemini models. The Claude search agent offers the highest Citation Precision, a score of and Claim Coverage, a score of . Meanwhile, OpenAI’s DeepResearch as well as the all other prior methods are unable to achieve a Citation Precision score beyond and a Claim Coverage score beyond . We also note that the human-written exemplars appear to exhibit rather low Citation Precision and Claim Coverage scores, however these scores are under-estimate the actual verifiability of human writing11footnotemark: 1. Overall, we see that prior LLM-based systems exhibit significant headroom for improvement.
5.1.2 DeepScholar-ref Provides a Strong Baseline for Generative Research Synthesis.
We compare the performance of DeepScholar-ref to OpenAI DeepResearch, search agents and open-source systems, finding that DeepScholar-ref provides a strong baseline with competitive performance against the other baselines across most metrics using the same or cheaper models. In comparison to OpenAI DeepResearch, DeepScholar-ref (GPT-4.1, o3) achieves a similar or higher Organization, Nugget Coverage, Relevance Rate, Reference Coverage, Citation Precision and Claim Coverage score. Notably, DeepScholar-ref achieves up to higher Verifiability scores but its Document Importance remain relatively low compared to OpenAI’s DeepResearch. In Appendix Table 6, we provide additional cost analysis, finding that DeepScholar-ref (GPT-4.1, o3) offers an efficient reference pipeline that is cheaper and faster than OpenAI DeepResearch.
Next, we compare the performance of DeepScholar-ref and search agents, finding DeepScholar-ref offers competitive and often stronger performance—specifically, averaged across the 5 baselines, using the same primary model, DeepScholar-ref increases Organization by , Nugget Coverage by , the Relevance Rate by , Reference Coverage by , Citation Precision by and Claim Coverage by . Lastly, we compare DeepScholar-ref (Llama-4) to the open-source research systems, all run with the Llama-4. We see that the prior open-source research systems exhibit trade-offs among the Knowledge Synthesis, Retrieval Quality and Verifiability dimensions. Compared to the best-performing prior open-source methods for each metric, DeepScholar-ref offers competitive Knowledge Synthesis performance, higher Relevance Rates, higher Reference Coverage, higher Citation Precision and higher Claim Coverage.
Overall, the strong relative performance of DeepScholar-ref likely reflects the efficiency of the data-processing semantic operators [37] that DeepScholar-ref uses to perform LLM-based filtering, ranking and summarization of sources to generate its report. Notably, DeepScholar-ref still demonstrates significant room for improvement and far from saturates DeepScholar-Bench, especially on key Knowledge Synthesis and Retrieval Quality metrics.
| Knowledge Synthesis | Retrieval Quality | Verifiability | Geo. Mean | |||||
| Org | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov () | ||
| DeepScholar-ref (GPT-4.1, Claude) | ||||||||
| arxiv.org Retrieval | .698 | .307 | .610 | .152 | .009 | .944 | .895 | .286 |
| parallel.ai Retrieval | .865 | .444 | .675 | .160 | .017 | .846 | .781 | .334 |
| taviliy.com Retrieval | .929∗ | .327 | .550 | .070 | .015 | .711 | .578 | .258 |
| Oracle Retrieval (arxiv.org) | .782 | .487 | .686 | 1.000∗ | 1.000 | .955 | .899∗ | .808 |
| Oracle Retrieval (All) | .778 | .528∗ | .680 | 1.000∗ | .822 | .941 | .828 | .782 |
| DeepScholar-ref (Llama-4) | ||||||||
| arxiv.org Retrieval | .206 | .241 | .436 | .103 | .008 | .674 | .851 | .195 |
| parallel.ai Retrieval | .246 | .265 | .559 | .114 | .015 | .223 | .543 | .186 |
| taviliy.com Retrieval | .111 | .229 | .532 | .030 | .016 | .442 | .676 | .153 |
| Oracle Retrieval (arxiv.org) | .202 | .316 | .681 | 1.000∗ | 1.000 | .658 | .868 | .590 |
| Oracle Retrieval (All) | .198 | .350 | .693 | 1.000∗ | .822 | .796∗ | .890 | .600 |
5.2 Understanding Opportunities for Improvement.
To analyze performance and opportunities for improvement, we conduct an ablation study, testing different retrievers, including two oracle settings. Table 3 shows the performance of DeepScholar-ref (GPT-4.1, Claude) and DeepScholar-ref (Llama-4), each with 3 different retrieval APIs: arxiv.org, the default in our main results, parallel.ai and tavily.com. In addition, our two oracle retrievers include the Oracle Retrieval (arxiv.org) setting and the Oracle Retrieval (All) setting, which provide the system with ArXiv references and all references, respectively, from the set of important references cited in exemplars, following our methodology for evaluating Reference Coverage.
Overall, the results demonstrate that key opportunities for improvement lie in both retrieval and knowledge synthesis capabilities. First, we see that DeepScholar-ref (GPT-4.1, Claude) with either oracle retriever nearly saturates performance on Retrieval Quality and Verifiability metrics, whereas the same method, using the arxiv.org, parallel.ai or taviliy.com retrievers, score much lower. Specifically, significant performance gaps exist for Reference Coverage and Document Importance, demonstrating the system struggles to navigate the live web and recover a diverse set of key, notable sources. Additionally, we also see that oracle retrievers improve the Nugget Coverage of either DeepScholar-ref methods by up to compared to the arxiv.org, parallel.ai or tavily.ai retrievers. Yet, the oracle settings still far from saturate Nugget Coverage, highlighting the AI system still struggle to effectively surface important facts and insights, even with high-quality sources.
5.3 Human Evaluation
To assess whether LLM-based metrics reflect human expert judgments, we conduct a human evaluation with 11 annotators, all Computer Science PhD students from four research universities across North America. In total, we collect over 300 human annotations in order to validate the LLM-based judges introduced by our automated evaluation for assessing knowledge synthesis and retrieval quality. We provide further details of our setup in Appendix A.3.4.
| Organization | |||
| Human / LLM | Reference | Generated | Tie |
| Reference | 14.29% | 0% | 14.29% |
| Generated | 0% | 57.14% | 14.29% |
| Tie | 0% | 0% | 0% |
| Nugget Importance | ||
| Human / LLM | Vital | Okay |
| Vital | 58.33% | 8.33% |
| Okay | 8.33% | 25.00% |
| Irrelevant | 0% | 0% |
| Reference Importance | ||
| Human / LLM | Not Imp. | Imp. |
| Not Imp. | 40.2% | 9.8% |
| Imp. | 24.2% | 25.7% |
Agreement analysis.
Table 4 shows the results of human evaluation study as confusion matrices taken between the majority vote of human annotators and the LLM-judge. Overall, the results show the robustness of our LLM-judges based on their strong alignment with the expert human annotators. Specifically, we observe a agreement score for pairwise comparisons judging Organization, a agreement score for nugget labeling to compute Nugget Coverage, and a agreement score for labeling reference importance to compute Reference Coverage. Notably, each of these tasks require reasoning about complex academic literature and lengthy candidate related works sections. The observed agreement rates from our study provide promising results, validating the use of automated LLM-judges and metrics to assess complex generative research synthesis tasks.
For Organization, we see from the confusion matrix that the main point of human- and LLM-judges most often agree on pairwise comparisons, with strong disagreements (i.e., humans preferring the reference report and the LLM preferring the generated report or vice versa) are rare. Moreover, of the disagreements that occur between human- and LLM-judges, the LLM mis-judgments are relatively equally balanced between picking the Reference report and the Generated Report.
For Nugget Importance, we observe that in addition to the strong agreement rate observed from the confusion matrix, we also see that the human majority vote find all LLM-generated nuggets to be relevant, indicating hallucinations are rare. Moreover we see that the false postive and false negative rate of the LLM-judge are similar, and both rather small, less than , once again indicating that severe LLM mis-labeling is rather rare.
Finally, for Reference Importance we observe an overall agreement score of , and importantly false negative rate of , i.e., when the LLM incorrectly labels a reference as important. The low false negative rate indicates a low likelihood of our Reference Coverage metric falsely penalizing systems. The larger off-diagonal mass (24.2%) reflects under-labeling of essential references by the LLM, suggesting that our Reference Coverage scores provide a rather conservative metric, measuring "recall" of only a subset of all truly important references for each query.
6 Related Work
Long-form Synthesis Benchmarks. While our work proposes a continually-updated, live benchmark using an automated data pipeline, several prior works instead provide expert-curated datasets for long-form research synthesis tasks, including ScholarQABench [6], OpenResearcher [66], DeepConsult [62], ResearcherBench [60], DeepResearch Bench [8], Deep Research Bench [10], SurGE [45], and LiveDRBench [16]. Unfortunately, these expert-curated benchmarks, are expensive to construct and update, can quickly become outdated, as new information becomes available, and risk data contamination, as new models are trained on publicly available data.
Alternatively, several recent benchmarks, including AcademicEval [63], LongBench-Cite [64] and SciIG [12], evaluate long-form generation tasks that do not require search over the live web, which is a key component of generative research synthesis and our benchmark. Other benchmarks focus on other long-form generation tasks, such as Wikipedia-like article generation [43], which differs substantially from our focus on complex research synthesis tasks. Crucially, unlike each of these prior works, our work proposes a live, continually-updated benchmark for evaluating generative research synthesis.
Factuality and Question Answering Benchmarks. While this work proposes a framework for studying complex, long-form research synthesis tasks, which lack an absolute notion of correctness and admit many possible reasonable answers, several recent works focus their evaluation on question answering (QA) and factuality benchmarks with short-form, easily verifiable answers. These prior benchmarks include SimpleQA [55], FRAMES [20], GAIA [31], BrowserComp [56], BrowserComp-Plus [7] WebWalkerQA [58], DeepResearch Arena [54] and others traditionally used to evaluate retrieval-augmented generation [53, 18, 61, 19, 21, 15, 49, 23]. Additionally, several benchmark develop automated dataset curation pipelines for live benchmark; however, their task focuses on short-form question-answering, as opposed to long-form report generation [36, 29, 17].
7 Conclusion
In this work, we introduced DeepScholar-bench, a live dataset and holistic, automated evaluation framework designed to rigorously benchmark an emerging class of systems designed for generative research synthesis. By automatically sourcing queries from high-quality, recent ArXiv papers, our benchmark mitigates the risks of data staleness and training contamination, while offering a real research synthesis task. Moreover, DeepScholar-bench provides an automated evaluation to holistically measure three critical dimensions: retrieval quality, knowledge synthesis and verifiability. We further release DeepScholar-ref, a reference pipeline, which we find provides a strong baseline for generative research synthesis. Overall our systematic evaluation of prior open-source systems, search agents, OpenAI’s DeepResearch and DeepScholar-ref demonstrates significant opportunities for future work, with no system surpassing a geometric mean of across all metrics. These results demonstrate both the difficulty of DeepScholar-bench and the exciting opportunity for further advancement in this space. We hope that DeepScholar-bench and DeepScholar-ref will support the development of more capable AI systems for generative research synthesis.
Acknowledgments
This research was supported in part by affiliate members and other supporters of the Stanford DAWN project, including Meta, Google, and VMware, as well as Cisco, SAP, and a Sloan Fellowship. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
Appendix A Appendix
A.1 DeepScholar-bench Dataset
We provide a detailed overview and schema of the DeepScholar-bench dataset in Figure 3. In addition, Table 5 includes the papers that are included in DeepScholar-June-2025.
| # | ArXiv ID | Title |
| 0 | 2506.02838v1 | TaxAgent: How Large Language Model Designs Fiscal Policy |
| 1 | 2506.02634v1 | KVCache Cache in the Wild: Characterizing and Optimizing KVCache Cache at a Large Cloud Provider |
| 2 | 2506.00958v1 | Speaking Beyond Language: A Large-Scale Multimodal Dataset for Learning Nonverbal Cues from Video-Grounded Dialogues |
| 3 | 2506.00832v1 | Counterfactual Activation Editing for Post-hoc Prosody and Mispronunciation Correction in TTS Models |
| 4 | 2506.00418v1 | Dual Debiasing for Noisy In-Context Learning for Text Generation |
| 5 | 2505.24754v1 | Don’t Reinvent the Wheel: Efficient Instruction-Following Text Embedding based on Guided Space Transformation |
| 6 | 2505.24575v1 | NexusSum: Hierarchical LLM Agents for Long-Form Narrative Summarization |
| 7 | 2506.00085v1 | COSMIC: Generalized Refusal Direction Identification in LLM Activations |
| 8 | 2505.23996v1 | Is Your Model Fairly Certain? Uncertainty-Aware Fairness Evaluation for LLMs |
| 9 | 2505.23353v1 | Synthetic Generation and Latent Projection Denoising of Rim Lesions in Multiple Sclerosis |
| 10 | 2505.22757v1 | Pre-Training Curriculum for Multi-Token Prediction in Language Models |
| 11 | 2506.02853v1 | Learning Pyramid-structured Long-range Dependencies for 3D Human Pose Estimation |
| 12 | 2506.02547v1 | Probabilistic Online Event Downsampling |
| 13 | 2506.01071v1 | Aligned Contrastive Loss for Long-Tailed Recognition |
| 14 | 2506.01037v1 | Self-supervised ControlNet with Spatio-Temporal Mamba for Real-world Video Super-resolution |
| 15 | 2506.00434v1 | Efficient 3D Brain Tumor Segmentation with Axial-Coronal-Sagittal Embedding |
| 16 | 2506.00333v1 | Test-time Vocabulary Adaptation for Language-driven Object Detection |
| 17 | 2505.24443v1 | Diversify and Conquer: Open-set Disagreement for Robust Semi-supervised Learning with Outliers |
| 18 | 2505.24334v1 | KairosAD: A SAM-Based Model for Industrial Anomaly Detection on Embedded Devices |
| 19 | 2505.23290v1 | Wav2Sem: Plug-and-Play Audio Semantic Decoupling for 3D Speech-Driven Facial Animation |
| 20 | 2505.23180v1 | Proximal Algorithm Unrolling: Flexible and Efficient Reconstruction Networks for Single-Pixel Imaging |
| 21 | 2505.22616v1 | PS4PRO: Pixel-to-pixel Supervision for Photorealistic Rendering and Optimization |
| 22 | 2505.22458v1 | Universal Domain Adaptation for Semantic Segmentation |
| 23 | 2505.22427v1 | RC-AutoCalib: An End-to-End Radar-Camera Automatic Calibration Network |
| 24 | 2505.22167v1 | Q-VDiT: Towards Accurate Quantization and Distillation of Video-Generation Diffusion Transformers |
| 25 | 2505.22552v1 | ClaimPKG: Enhancing Claim Verification via Pseudo-Subgraph Generation with Lightweight Specialized LLM |
| 26 | 2504.21752v1 | VDDP: Verifiable Distributed Differential Privacy under the Client-Server-Verifier Setup |
| 27 | 2504.21282v1 | Birdie: Natural Language-Driven Table Discovery Using Differentiable Search Index |
| 28 | 2504.17448v1 | CHASe: Client Heterogeneity-Aware Data Selection for Effective Federated Active Learning |
| 29 | 2504.14861v1 | Stitching Inner Product and Euclidean Metrics for Topology-aware Maximum Inner Product Search |
| 30 | 2504.06975v1 | AWDIT: An Optimal Weak Database Isolation Tester |
| 31 | 2506.01833v1 | SPACE: Your Genomic Profile Predictor is a Powerful DNA Foundation Model |
| 32 | 2506.00382v1 | Spectral Insights into Data-Oblivious Critical Layers in Large Language Models |
| 33 | 2506.00205v1 | Unlocking the Power of Rehearsal in Continual Learning: A Theoretical Perspective |
| 34 | 2505.24835v1 | Timing is important: Risk-aware Fund Allocation based on Time-Series Forecasting |
| 35 | 2505.24203v1 | Aligning Protein Conformation Ensemble Generation with Physical Feedback |
| 36 | 2506.02847v1 | CLONE: Customizing LLMs for Efficient Latency-Aware Inference at the Edge |
| 37 | 2505.22194v1 | Refining Datapath for Microscaling ViTs |
| 38 | 2505.11554v1 | Multi-Objective Memory Bandwidth Regulation and Cache Partitioning for Multicore Real-Time Systems |
| 39 | 2505.08071v1 | NMP-PaK: Near-Memory Processing Acceleration of Scalable De Novo Genome Assembly |
| 40 | 2504.06211v1 | Need for zkSpeed: Accelerating HyperPlonk for Zero-Knowledge Proofs |
| 41 | 2504.19283v1 | Efficient Serverless Cold Start: Reducing Library Loading Overhead by Profile-guided Optimization |
| 42 | 2504.11007v1 | Kubernetes in the Cloud vs. Bare Metal: A Comparative Study of Network Costs |
| 43 | 2504.09307v1 | Lumos: Efficient Performance Modeling and Estimation for Large-scale LLM Training |
| 44 | 2506.02750v1 | Learning Binarized Representations with Pseudo-positive Sample Enhancement for Efficient Graph Collaborative Filtering |
| 45 | 2505.23452v1 | What About Emotions? Guiding Fine-Grained Emotion Extraction from Mobile App Reviews |
| 46 | 2505.21811v1 | Revisiting Self-attention for Cross-domain Sequential Recommendation |
| 47 | 2505.20227v1 | Measure Domain’s Gap: A Similar Domain Selection Principle for Multi-Domain Recommendation |
| 48 | 2505.19356v1 | Optimized Text Embedding Models and Benchmarks for Amharic Passage Retrieval |
| 49 | 2505.19307v1 | Aligning Web Query Generation with Ranking Objectives via Direct Preference Optimization |
| 50 | 2505.17507v1 | Benchmarking Recommendation, Classification, and Tracing Based on Hugging Face Knowledge Graph |
| 51 | 2505.12791v1 | Unlearning for Federated Online Learning to Rank: A Reproducibility Study |
| 52 | 2505.07166v1 | Pre-training vs. Fine-tuning: A Reproducibility Study on Dense Retrieval Knowledge Acquisition |
| 53 | 2505.03484v1 | STAR-Rec: Making Peace with Length Variance and Pattern Diversity in Sequential Recommendation |
| 54 | 2505.00552v1 | Graph Spectral Filtering with Chebyshev Interpolation for Recommendation |
| 55 | 2504.20458v1 | Search-Based Interaction For Conversation Recommendation via Generative Reward Model Based Simulated User |
| 56 | 2504.18383v1 | Bridge the Domains: Large Language Models Enhanced Cross-domain Sequential Recommendation |
| 57 | 2504.17519v1 | Replication and Exploration of Generative Retrieval over Dynamic Corpora |
| 58 | 2504.15849v1 | NLCTables: A Dataset for Marrying Natural Language Conditions with Table Discovery |
| 59 | 2504.14991v1 | Understanding Accuracy-Fairness Trade-offs in Re-ranking through Elasticity in Economics |
| 60 | 2504.14243v1 | Unconstrained Monotonic Calibration of Predictions in Deep Ranking Systems |
| 61 | 2504.12900v1 | FashionDPO:Fine-tune Fashion Outfit Generation Model using Direct Preference Optimization |
| 62 | 2504.09935v1 | Constrained Auto-Regressive Decoding Constrains Generative Retrieval |
A.2 Overview of Baselines and Experimental Setup
We briefly overview all of the baseline systems we evaluate, which includes recent state-of-the-art generative research systems as well as DeepScholar-ref on DeepScholar-bench. We benchmark open-source research systems, including DeepResearcher [65], STORM [43] and OpenScholar [6], search agents, with Llama-4-Scout-17B-16E-Instruct [30], GPT-4.1-2025-04-14 [34], o3-2025-04-16 [34], Claude-opus-4-20250514 [3], and Gemini-2.5-pro [14] models, OpenAI’s o3-deep-research [34], and DeepScholar-ref.